three times the largest of three consecutive even integers is twenty-two more than the sum of the other two. Find the integers

Answer by checkley77(12844)

Let x, x+2 & x=4 be the 3 integers.
3(x+4)=x+x+2
3x+12=2x+2
3x-2x=2-12
x=-10 ans. for the first integer.
-10+2=-8 ans. for the middle integer.
-10=4=-6 ans. for the largest integer.
Proof;
3*-6=-10-8
-18=-18